# Akide Liu1  Jing Liu1  Zizheng Pan1  Yefei He2 Gholamreza Haffari1Bohan Zhuang1, 2

### v1

@@ 4.3 Efficiency Discussion
Restoration efficiency. We then analyze the additional memory cost incurred during the restoration process, which During the magnitude rescaling phase, we save an additional norm vector for the corresponding layers in the KV cache. It is important to note that the norm vector is in the shape of $\mathbb{R}^{b\times s\times 1}$, which has a single channel dimension compared to the fully ranked original KV states. Additionally, we suggest that the retention threshold can be set to 0.05. Therefore, we have $brh(0.05(s+n))$ tokens retained without compression. Finally, our overall memory requirement is given by $(3.1h+2)br(s+n)$. The detailed derivation is shown in the Appendix E.
@@ 5 Experiments
Implementation details. Our experiments are based on representative model families of LLMs, including a compact LLM Phi-3-Mini [23] and an MoE LLM Mixtral-8x7B [22]. Additionally, we adopt LLaMA-3 [6] 8B and 70B models to explore how our method generalizes to larger LLMs. We sample ten tasks from lm-eval-harness [32], including COPA [24], MathQA [25], OpenBookQA [26], PiQA [27], RTE [28], Winogrande [29], XSUM [30], and CNN/Daily Mail [31]. We also evaluate long-sequence generation on LongBench [33]. We compare ours with fully cached baseline, and other methods such as round-to-nearest quantization (RTN) [73], SmoothQuant [70] and KIVI [11].
@@
LongBench. We also conducted experiments to evaluate performance and quality in long-sequence generation using the LongBench dataset [33], shown in Table 1. Our experiments applied MiniCache over several models: LLaMA-2-7B-Chat, LLaMA-2-13B-Chat, Mistral-7B, and Mistral-7B-Instruct. It is important to note that our MiniCache method maintains orthogonality with all existing quantization and sparsity (refers to Table A) methods at both the model and token-wise levels. When combined with KIVI-4bit KV cache quantization, our approach achieved a compression rate of 5.02×5.02$\times$5.02 ×, with minimal impact on accuracy across various challenging long-context generation tasks. The combination of MiniCache and KIVI-4bit KV cache quantization demonstrates significant memory savings without compromising the model’s ability to handle long sequences effectively. This highlights the potential of our method to optimize large language models for tasks requiring extensive context, making them more efficient and scalable for real-world applications.
@@ 6 Ablation Study
Table 2: Comparison of token retention threshold $\gamma$ by LLaMA-2-7B [5] on three benchmarks.
The effect of interpretation parameter $t$. In our exploration of the effects of the interpretation parameter $t$ on performance, particularly in relation to the relative magnitude ratio of adjacent layers, as shown in Figure 6. We maintain all settings constant, starting from layer $S=16$ (halfway through the layers of LLaMA-3-8B), and vary the interpretation parameter $t$ from 0.3 to 0.7. Our findings reveal several key points: When $t=0.5$, the process resembles average merging, which is less effective for cross-layer merging. In contrast, when $t=0.6$ is optimal, the merged representation exhibits the most robust performance, while indicating that more information is derived from the second term (${\bm{x}}^{l}$) of the SLERP.
@@ 7 Conclusion and Future Work
Appendix A Additional Experiments Results
Comparison with token sparsity methods. We also compare MiniCache with the sparsity-based method H2O [14]. Our method outperforms H2O in most tasks on LongBench. Notably, our approach addresses inter-layer redundancy, whereas H2O focuses on intra-layer redundancy. Our approaches are orthogonal to sparsity-based methods.
Table A: Compare MiniCache with token sparsity based method H2O, using mistral-7B-instruct on LongBench dataset.
@@ Appendix A Additional Experiments Results
Appendix B Additional Related Works
Preliminaries of KV cache in LLMs. LLM inference has been significantly improved in recent works [9, 74, 13] by optimizing the KV cache management. Overall, this line of research is typically done in two steps: 1) First, at the prefilling stage, LLM calculates the initial KV caches at each attention layer based on the input sequence and decodes the first token. Second, 2) at the decoding Stage, LLM autoregressively predicts the next token, where its KV cache at each attention layer is added to the overall caches. While existing works have compressed KV cache in different aspects (e.g., quantization [11, 12], token pruning [14, 16] ).
KV cache compression. In the prior study, various strategies for enhancing efficient transformer architectures are discussed, covering a spectrum of techniques aimed at optimizing performance and managing resource constraints. These methods include attention optimization [46, 47, 76], grouping queries [43, 42], sparse KV caching [16, 77, 78], shrinking tokens [79, 15], and improving long-context generation. Significant contributions come from projects such as H2O [15], GEAR [15], and KIVI [11]. Additionally, efforts to alleviate KV cache bottlenecks include strategies like multi-query attention [42] and multi-group attention [43], which propose reducing the number of heads in the KV cache. However, these methods often necessitate retraining or fine-tuning models from scratch. Other approaches focus on diminishing the size of the KV cache by selectively evicting less important tokens [14] and enhancing the system architecture by offloading the KV cache [80] or integrating techniques such as virtual memory and paging [81] into the attention mechanism.
@@ Appendix C Discussions and Limitations
Alternative merging function. During our preliminary exploration, we first consider another simpler merge function maximum norm-preserving interpolation for cross-layer compression. This function is designed to maintain the maximum norm of the vectors involved, ensuring that the most significant features are preserved during the merging process. The maximum norm-preserving interpolation in terms of $F_{\mathrm{max}}$ can be defined as follows:
$F_{\mathrm{max}}({\bm{x}}^{l},{{\bm{x}}}^{l-1})=\frac{\bar{{\bm{x}}}^{l,l-1}}{\lVert\bar{{\bm{x}}}^{l,l-1}\rVert}\cdot\mathrm{Max}(\lVert{\bm{x}}^{l}\rVert,\lVert{\bm{x}}^{l-1}\rVert)$ (A)
@@
Societal impact. Our work represents a preliminary exploration of KV cache Compression in the depth dimension, an area that has been relatively unexplored and yet presents a significant bottleneck in large language models (LLMs). The proposed MiniCache provides a solution to improve the efficiency of LLM generation and is adaptable to existing intra-layer-based KV cache pruning and quantization technologies. Additionally, we proposed a simple yet effective approach that merges similar KV cache states in a cross-layer manner and effectively restores performance through a novel restoration technique. Our observations indicate that cross-layer similarity and mergeability can be applied in broad-efficient applications for post-training optimization in low-resource scenarios, such as deployment on mobile devices. Moreover, with effective optimization of KV cache memory demand, MiniCache further enhances long-context generation, which is a crucial paradigm for real-world applications, such as understanding concepts in textbooks. We hope our work supports the two key challenges of batch inference and long-context generation in the LLM industry and further research.
Furthermore, in addition to the significant contributions of MiniCache for KV cache Compression, several challenges remain that are common to large language models (LLMs). Issues such as the truthfulness and security of LLMs are still unresolved. Ensuring the accuracy and reliability of generated content is critical, as LLMs can sometimes produce plausible but incorrect or misleading information. Additionally, safeguarding against security vulnerabilities, such as adversarial attacks or data leakage, is paramount to maintaining the integrity and confidentiality of user interactions. Addressing these challenges requires ongoing research and development to enhance the robustness and trustworthiness of LLMs, alongside the advancements in computational efficiency and performance provided by innovations like MiniCache.
@@ Appendix D Additional Implementation Details
Cross-Layer merging and restoration algorithm. As outlined in Alg. 2, involves several crucial steps to ensure efficient memory usage while maintaining the integrity of the Key-Value (KV) Cache states. Initially, given the KV cache (${\bm{E}}_{k,v}^{l}$), norm values ($\lVert{\bm{X}}_{k,v}^{l}\rVert$), unmerged tokens (${\bm{R}}_{k,v}^{l}$), retention indices (${\mathbb{I}}_{k,v}$), and the next tokens (${\bm{t}}^{l}$, ${\bm{t}}^{l-1}$), the algorithm proceeds by rescaling the magnitude of the KV pairs. Specifically, $\hat{{\bm{X}}}_{k}^{l}$ and $\hat{{\bm{X}}}_{v}^{l}$ are computed by multiplying the normalized KV pairs ${\bm{E}}_{k,v}^{l}$ with their respective magnitude norms $\lVert{\bm{X}}_{k,v}^{l}\rVert$. Following this, the algorithm restores unmerged tokens using the retention indices, updating $\hat{{\bm{X}}}_{k}^{l}$ and $\hat{{\bm{X}}}_{v}^{l}$ accordingly. Next, the new tokens ${\bm{t}}_{k}$ and ${\bm{t}}_{v}$ are concatenated to the rescaled KV pairs along the token dimension. This augmented KV cache undergoes a softmax attention mechanism where the attention scores ${\bm{A}}$ are computed by taking the dot product of the query token $\bm{t}_{q}$ with the transposed keys $(\hat{{\bm{X}}}_{k}^{l})^{\top}$. The output token ${\bm{t}}_{O}$ is then obtained by multiplying the attention scores ${\bm{A}}$ with the values $\hat{{\bm{X}}}_{v}^{l}$. In cases where the previous token ${\bm{t}}^{l-1}$ exists, the algorithm performs a compression step. It concatenates the existing KV cache ${\bm{E}}_{k,v}^{l}$ with the merged tokens resulting from the current and previous layers, effectively reducing redundancy and optimizing memory. If ${\bm{t}}^{l-1}$ is not available, the KV cache is updated by simply concatenating ${\bm{E}}_{k,v}^{l}$ with the new tokens ${\bm{t}}_{k,v}^{l}$, deferring the compression until the next iteration. The final output token ${\bm{t}}_{O}$ is then returned, concluding the decoding process. In the merging function, the algorithm normalizes the KV pairs from both the current and previous layers. It calculates the angular distance $\Omega$ between the normalized vectors, ensuring that the interpolation occurs along the shortest path on the unit sphere. The merged KV cache is then obtained by weighted interpolation of the normalized vectors, preserving the geometric and semantic integrity of the original states. This comprehensive process allows MiniCache to achieve substantial memory efficiencies while maintaining the functional characteristics of the KV pairs across transformer layers.
@@
MiniCache execution flow. Figure A delineates the pre-filling and decoding logic for the MiniCache framework, which incorporates cross-layer merging and error suppression to achieve memory efficiency and maintain functional integrity. Initially, in Step 1, the KV cache is fetched from the previous layer (Layer L-1) during the pre-filling phase. In Step 2, the fetched KV pairs from the current layer ($\chi^{L}$) are merged with the KV pairs from the preceding layer ($\chi^{L-1}$), reducing redundancy through a merge function. Subsequently, in Step 3, the merged KV pairs are cached for future use, representing a consolidated data set from multiple layers. During the decoding phase, Step 4 involves deleting unnecessary or redundant KV pairs to optimize memory usage. In Step 5, the decoder fetches the required KV pairs from the cache for output generation. Step 6 applies error suppression mechanisms to the fetched KV pairs, including rescaling and retention recovery, to minimize errors introduced during the merging and compression processes. Finally, in Step 7, the cache is updated with the final KV pairs post-error suppression and adjustments, ensuring the cache contains the most accurate and efficient representation of the KV pairs for subsequent layers. This comprehensive approach guarantees substantial memory efficiencies while preserving the critical functional characteristics of the original KV pairs across transformer layers.
@@ 16            Standard Attention & Standard Cache
       $d({\bm{X}}^{l},{\bm{X}}^{l-1})_{k,v}=\frac{1}{\pi}\cdot\Omega_{k,v}$ /* distance metrics */
6       ;
       ${\mathbb{I}}_{k,v}=\{i\mid d_{i}<d_{\text{min}}+(d_{\text{max}}-d_{\text{min}})\cdot\gamma\}$ /* retention indices */
7      ;
       ${\bm{R}}_{k}^{l,l-1}\leftarrow{\bm{X}}_{k}^{l,l-1}[{\mathbb{I}}_{k}],{\bm{R}}_{v}^{l,l-1}\leftarrow{\bm{X}}_{v}^{l,l-1}[{\mathbb{I}}_{v}]$ /* unmerged tokens */
8       ;
@@
10
11procedure MiniCache Decoding:
@@
17
18function MiniCache Merge(${\bm{X}}^{l},{\bm{X}}^{l-1},t$):
@@ Appendix E Detailed Efficiency Derivation
$4brh(s+n)\cdot\frac{1}{2}\cdot\frac{1}{2}=brh(s+n)$
@@ Appendix F Detailed Experiment Results
Table B: Detailed performance comparison on GSM8K dataset with LLaMA-3-70B
@@
Table C: Detailed performance comparison on COQA dataset with LLaMA-3-70B
@@
Table D: Detailed performance comparison on TruthfulQA dataset with LLaMA-3-70B
@@
Table E: Detailed performance comparison on GSM8K dataset with LLaMA-3-8B
@@
Table F: Detailed performance comparison on COQA dataset with LLaMA-3-8B
@@
Table G: Detailed performance comparison on TruthfulQA dataset with LLaMA-3-8B
@@
Table H: Detailed performance comparison on GSM8K dataset with Mixtral-8x7B
@@
Table I: Detailed performance comparison on COQA dataset with Mixtral-8x7B
@@
Table J: Detailed performance comparison on TruthfulQA dataset with Mixtral-8x7B
@@
Table K: Detailed performance comparison on GSM8K dataset with Phi-3-Mini
@@
Table L: Detailed performance comparison on COQA dataset with Phi-3-Mini
@@
Table M: Detailed performance comparison on TruthfulQA dataset with Phi-3-Mini

### v2

@@ 4.3 Efficiency Discussion
Restoration efficiency. We then analyze the additional memory cost incurred during the restoration process, which During the magnitude rescaling phase, we save an additional norm vector for the corresponding layers in the KV cache. It is important to note that the norm vector is in the shape of $\mathbb{R}^{b\times s\times 1}$, which has a single channel dimension compared to the fully ranked original KV states. Additionally, we suppose that the retention threshold can be set to 0.05. Therefore, we have $brh(0.05(s+n))$ tokens retained without compression. Finally, our overall memory requirement is given by $(3.1h+2)br(s+n)$. The detailed derivation is shown in the Appendix E.
@@ 5 Experiments
Implementation details. Our experiments are based on representative model families of LLMs, including a compact LLM Phi-3-Mini [23] and an MoE LLM Mixtral-8x7B [22]. Additionally, we adopt LLaMA-3 [6] 8B and 70B models to explore how our method generalizes to larger LLMs. We sample ten tasks from lm-eval-harness [32], including COPA [24], MathQA [25], OpenBookQA [26], PIQA [27], RTE [28], WinoGrande [29], XSUM [30], and CNN/Daily Mail [31]. We also evaluate long-sequence generation on LongBench [33]. We compare our method with a fully cached baseline, and other methods such as round-to-nearest quantization (RTN) [73], SmoothQuant [70] and KIVI [11].
@@
LongBench. We also conduct experiments to evaluate performance and quality in long-sequence generation using the LongBench dataset [33], as shown in Table 1. Our experiments applied MiniCache over several models: LLaMA-2-7B-Chat, LLaMA-2-13B-Chat, Mistral-7B, and Mistral-7B-Instruct. It is important to note that our MiniCache method maintains orthogonality with all existing quantization and sparsity (refers to Table A) methods at both the model and token-wise levels. When combined with KIVI-4bit KV cache quantization, our approach achieves a compression rate of 5.02×5.02$\times$5.02 ×, with minimal impact on accuracy across various challenging long-context generation tasks. The combination of MiniCache and KIVI-4bit KV cache quantization demonstrates significant memory savings without compromising the model’s ability to handle long sequences effectively. This highlights the potential of our method to optimize large language models for tasks requiring extensive context, making them more efficient and scalable for real-world applications.
@@ 6 Ablation Study
Table 2: Comparisons of various token retention thresholds $\gamma$ by LLaMA-2-7B [5] on three benchmarks.
@@
The effect of interpretation parameter $t$. We explore the effects of the interpretation parameter $t$ on performance, particularly in relation to the relative magnitude ratio of adjacent layers, as shown in Figure 6. We maintain all settings constant, starting from layer $S=16$ (halfway through the layers of LLaMA-3-8B), and vary the interpretation parameter $t$ from 0.3 to 0.7. Our findings reveal several key points. When $t=0.5$, the process resembles average merging, which is less effective for cross-layer merging. In contrast, when $t=0.6$ is optimal, the merged representation exhibits the most robust performance, while indicating that more information is derived from the second term (${\bm{x}}^{l}$) of the SLERP.
@@ 7 Conclusion and Future Work
Appendix A Additional Experiment Results
Comparisons with token sparsity methods. We also compare MiniCache with the sparsity-based method H2O [14]. Our method outperforms H2O in most tasks on LongBench. Notably, our approach focuses on reducing inter-layer redundancy, whereas H2O focuses on intra-layer redundancy. Our approaches are orthogonal to sparsity-based methods.
Table A: Comparison between MiniCache with token sparsity based method H2O, using mistral-7B-instruct on LongBench dataset.
@@ Appendix A Additional Experiment Results
Appendix B Additional Related Work
Preliminaries of KV cache in LLMs. LLM inference has been significantly improved in recent works [9, 74, 13] by optimizing the KV cache management. Overall, this line of research is typically done in two steps: 1) First, at the prefilling stage, LLM calculates the initial KV caches at each attention layer based on the input sequence and decodes the first token. Second, 2) at the decoding stage, LLM autoregressively predicts the next token, where its KV cache at each attention layer is added to the overall caches. Existing works have compressed KV cache in different aspects (e.g., quantization [11, 12], token pruning [14, 16] ).
KV cache compression. In the prior study, various strategies for enhancing efficient transformer architectures are discussed, covering a spectrum of techniques aimed at optimizing performance and managing resource constraints. These methods include attention optimization [46, 47, 76], grouping queries [43, 42], sparse KV caching [16, 77, 78], shrinking tokens [79, 15], and improving long-context generation. Significant contributions come from projects such as H2O [15], GEAR [15], and KIVI [11]. Additionally, efforts to alleviate KV cache bottlenecks include strategies like multi-query attention [42] and multi-group attention [43], which propose reducing the number of heads in the KV cache. However, these methods often necessitate retraining or fine-tuning models. Other approaches focus on diminishing the size of the KV cache by selectively evicting less important tokens [14] and enhancing the system architecture through technologies like offloading the KV cache [80] or integrating techniques such as virtual memory and paging [81] into the attention mechanism.
@@ Appendix C Discussions and Limitations
Alternative merging function. During our preliminary exploration, we initially considered an alternative, simpler merge function for cross-layer compression: maximum norm-preserving interpolation. This function is designed to maintain the maximum norm of the vectors involved, ensuring that the most significant features are preserved during the merging process. The maximum norm-preserving interpolation in terms of $F_{\mathrm{max}}$ can be defined as follows:
$F_{\mathrm{max}}({\bm{x}}^{l},{{\bm{x}}}^{l-1})=\frac{\bar{{\bm{x}}}^{l,l-1}}{\lVert\bar{{\bm{x}}}^{l,l-1}\rVert}\cdot\mathrm{Max}(\lVert{\bm{x}}^{l}\rVert,\lVert{\bm{x}}^{l-1}\rVert).$ (A)
@@
Societal impact. Our work shows a preliminary exploration of KV cache Compression in the depth dimension, a relatively unexplored yet critically bottlenecked area in large language models (LLMs). The proposed MiniCache provides a solution to improve the efficiency of LLM generation and is adaptable to existing intra-layer-based KV cache pruning and quantization technologies. Additionally, we proposed a simple yet effective approach that merges similar KV cache states in a cross-layer manner and effectively restores performance through a novel restoration technique. Our observations indicate that cross-layer similarity and mergeability can be applied in broad-efficient applications for post-training optimization in low-resource scenarios, such as deployment on mobile devices. Moreover, with effective optimization of KV cache memory demand, MiniCache further enhances long-context generation, which is a crucial paradigm for real-world applications, such as understanding concepts in textbooks. We aim for our work to advance the boundaries of two key challenges in the LLM industry and research: batch inference and long-context generation.
Furthermore, in addition to the significant contributions of MiniCache for KV cache Compression, several challenges remain that are common to LLMs. Issues such as the truthfulness and security of LLMs are still unresolved. Ensuring the accuracy and reliability of generated content is critical, as LLMs can sometimes produce plausible but incorrect or misleading information. Additionally, safeguarding against security vulnerabilities, such as adversarial attacks or data leakage, is paramount to maintaining the integrity and confidentiality of user interactions. Addressing these challenges requires ongoing research and development to enhance the robustness and trustworthiness of LLMs. This effort must proceed alongside advancements in computational efficiency and performance, as exemplified by innovations like MiniCache.
@@ Appendix D Additional Implementation Details
Cross-Layer merging and restoration algorithm. As outlined in Alg. 2, MiniCache algorithm involves several crucial steps to ensure efficient memory usage while maintaining the integrity of the Key-Value (KV) Cache states. Initially, given the KV cache ${\bm{E}}_{k,v}^{l}$, norm values $\lVert{\bm{X}}_{k,v}^{l}\rVert$ unmerged tokens ${\bm{R}}_{k,v}^{l}$, retention indices ${\mathbb{I}}_{k,v}$, and the next tokens ${\bm{t}}^{l}$, ${\bm{t}}^{l-1}$, the algorithm proceeds by rescaling the magnitude of the KV pairs. Specifically, $\hat{{\bm{X}}}_{k}^{l}$ and $\hat{{\bm{X}}}_{v}^{l}$ are computed by multiplying the normalized KV pairs ${\bm{E}}_{k,v}^{l}$ with their respective magnitude norms $\lVert{\bm{X}}_{k,v}^{l}\rVert$. Following this, the algorithm restores unmerged tokens using the retention indices, updating $\hat{{\bm{X}}}_{k}^{l}$ and $\hat{{\bm{X}}}_{v}^{l}$ accordingly. Next, the new tokens ${\bm{t}}_{k}$ and ${\bm{t}}_{v}$ are concatenated to the rescaled KV pairs along the token dimension. This augmented KV cache undergoes a softmax attention mechanism where the attention scores ${\bm{A}}$ are computed by taking the dot product of the query token $\bm{t}_{q}$ with the transposed keys $(\hat{{\bm{X}}}_{k}^{l})^{\top}$. The output token ${\bm{t}}_{O}$ is then obtained by multiplying the attention scores ${\bm{A}}$ with the values $\hat{{\bm{X}}}_{v}^{l}$. In cases where the previous token ${\bm{t}}^{l-1}$ exists, the algorithm performs a compression step. It concatenates the existing KV cache ${\bm{E}}_{k,v}^{l}$ with the merged tokens resulting from the current and previous layers, effectively reducing redundancy and optimizing memory. If ${\bm{t}}^{l-1}$ is not available, the KV cache is updated by simply concatenating ${\bm{E}}_{k,v}^{l}$ with the new tokens ${\bm{t}}_{k,v}^{l}$, deferring the compression until the next iteration. The final output token ${\bm{t}}_{O}$ is then returned, concluding the decoding process. In the merging function, the algorithm normalizes the KV pairs from both the current and previous layers. It calculates the angular distance $\Omega$ between the normalized vectors, ensuring that the interpolation occurs along the shortest path on the unit sphere. The merged KV cache is then obtained by weighted interpolation of the normalized vectors, preserving the geometric and semantic integrity of the original states. This comprehensive process allows MiniCache to achieve substantial memory efficiencies while maintaining the functional characteristics of the KV pairs across transformer layers.
@@
MiniCache execution flow. Figure A delineates the pre-filling and decoding logic for the MiniCache framework, which incorporates cross-layer merging and error suppression to achieve memory efficiency and maintain functional integrity. Initially, in Step 1, the KV cache is fetched from the previous layer (Layer $L-1$) during the pre-filling phase. In Step 2, the fetched KV pairs from the current layer $\chi^{L}$ are merged with the KV pairs from the preceding layer $\chi^{L-1}$, reducing redundancy through a merge function. Subsequently, in Step 3, the merged KV pairs are cached for future use, representing a consolidated data set from multiple layers. During the decoding phase, Step 4 involves deleting unnecessary or redundant KV pairs to optimize memory usage. In Step 5, the decoder fetches the required KV pairs from the cache for output generation. Step 6 applies error suppression mechanisms to the fetched KV pairs, including rescaling and retention recovery, to minimize errors introduced during the merging and compression processes. Finally, in Step 7, the cache is updated with the final KV pairs post-error suppression and adjustments, ensuring the cache contains the most accurate and efficient representation of the KV pairs for subsequent layers. This comprehensive approach guarantees substantial memory efficiencies while preserving the critical functional characteristics of the original KV pairs across transformer layers.
@@ 16            Standard Attention & Standard Cache
       $d({\bm{X}}^{l},{\bm{X}}^{l-1})_{k,v}=\frac{1}{\pi}\cdot\Omega_{k,v}$ // distance metrics
        ${\mathbb{I}}_{k,v}=\{i\mid d_{i}<d_{\text{min}}+(d_{\text{max}}-d_{\text{min}})\cdot\gamma\}$ // retention indices
       ${\bm{R}}_{k}^{l,l-1}\leftarrow{\bm{X}}_{k}^{l,l-1}[{\mathbb{I}}_{k}],{\bm{R}}_{v}^{l,l-1}\leftarrow{\bm{X}}_{v}^{l,l-1}[{\mathbb{I}}_{v}]$ // unmerged tokens
@@
7
8procedure MiniCache Decoding:
@@
14
15function MiniCache Merge(${\bm{X}}^{l},{\bm{X}}^{l-1},t$):
@@ Appendix E Detailed Efficiency Derivation
$4brh(s+n)\cdot\frac{1}{2}\cdot\frac{1}{2}=brh(s+n).$
@@ Appendix F Detailed Experiment Results
Table B: Detailed performance comparison on GSM8K dataset with LLaMA-3-70B.
@@
Table C: Detailed performance comparison on COQA dataset with LLaMA-3-70B.
@@
Table D: Detailed performance comparison on TruthfulQA dataset with LLaMA-3-70B.
@@
Table E: Detailed performance comparison on GSM8K dataset with LLaMA-3-8B.
@@
Table F: Detailed performance comparison on COQA dataset with LLaMA-3-8B.
@@
Table G: Detailed performance comparison on TruthfulQA dataset with LLaMA-3-8B.
@@
Table H: Detailed performance comparison on GSM8K dataset with Mixtral-8x7B.
@@
Table I: Detailed performance comparison on COQA dataset with Mixtral-8x7B.
@@
Table J: Detailed performance comparison on TruthfulQA dataset with Mixtral-8x7B.
@@
Table K: Detailed performance comparison on GSM8K dataset with Phi-3-Mini.
@@
Table L: Detailed performance comparison on COQA dataset with Phi-3-Mini.
@@
Table M: Detailed performance comparison on TruthfulQA dataset with Phi-3-Mini.